Please I need help with this problem, please show work.
7q^-4•q^5r^-5÷-14q^-5r^-6

Answer by jessica43(140)

To solve this problem you are going to use two of the basic rules of exponents: 1. Whenever you multiply two terms with the same base, you can add the exponents:
(x^m)*(x^n) = x^(m + n)
2. Whenever you divide two terms with the same base, you can subtract the exponents:
(x^m)/(x^n) = x^(m-n)
Now we will use these rules to solve this problem:
(7(q^-4)(q^5)(r^-5))/((-14)(q^-5)(r^-6))
=(7(q^1)(r^-5))/((-14)(q^-5)(r^-6)) (rule 1: (q^-4)*(q^5)= q^(-4+5) = q^1)
=(7(q^6)(r^-5))/((-14)(r^-6)) (rule 2: (q^1)/(q^-5)= q^(1-(-5)= q^6)
=(7(q^6)(r^1))/(-14) (rule 2: (r^-5)/(r^-6) = r^(-5-(-6)) = r^1)
=((q^6)(r^1))/(-2) (simplified: -7/14 = -1/2)
So your simplified answer is:
-(((q^6)(r))/2)
Hope this helps!